A divalent element x has a specific heat of 0.106 cal/g/°C. When a solution of the sulphate of x is electrolysed, 4.85g of x is deposited at the cathode and one litre of oxygen gas, measured at 16°C and 743mmHg pressure is evolved at the anode. Calculate the relative atomic mass of x (molar heat capacity =6.4 cal/mol/°C)
1
Expert's answer
2020-08-22T08:25:14-0400

Ideal gas low:

PV=nRT

Molar gas constant

R=0.082058 L×atm/(K×mol)

T = 16 + 273.15 = 289.15 K

P = 0.977 atm

n(O2) = PV/RT

n(O2) = 0.977×1/0.082058×289.15 = 0.041 mol

Electrolysis reaction:

2XSO4 + 2H2O → 2X + 2H2SO4 + O2

n(X) = 2n(O2)

n(X) = 2×0.041 = 0.082 mol

M = m/n

M(X) = 4.85/0.082 = 59.14 g/mol

Answer: 59.14 g/mol